### 24 Mar 2026 v1

# Piecewise M-Stationarity and Related Algorithms for Mathematical Programs with Complementarity Constraints

Abstract
This study explores B-stationarity of mathematical programs with complementarity constraints (MPCCs) and convergence behavior of MPCC algorithms. Special attention is given to the cases with biactive complementarity constraints. First, we propose the concept of piecewise M-stationarity and prove its equivalence to B-stationarity under MPCC-ACQ. Then, we investigate convergence properties of the NCP-based bounding methods we proposed in [31], without requiring MPCC-LICQ; an interpretation of the algorithm’s behavior together with the concept of piecewise M-stationarity leads to a cost reduction in B-stationarity verification. In addition, practical issues related to convergence to non-strongly stationary solutions are discussed, which shows that the NCP-based complementarity reformulations have an advantage in avoiding unbounded multipliers near these solutions.
Keywords: Complementarity constraints; MPCCs; B-stationarity; Stationarity conditions; Constraint qualifications; Smoothed NCP functions
1 Introduction
We consider mathematical programs with complementarity constraints (MPCCs) of the form
$\min$ $f(z)$ (1)
${\rm s.t.}$ $g(z)\leq 0,$
$h(z)=0,$
$0\leq G(z)\perp H(z)\geq 0,$
where $(f,g,h,G,H):\mathbb{R}^{n}\to\mathbb{R}^{1+n_{g}+n_{h}+m+m}$ are differentiable functions. At a feasible point $\bar{z}$ of the MPCC, define the following index sets:
$I_{g}(\bar{z})$ $=\{i\,|\,g_{i}(\bar{z})=0\},$ (2)
$I_{h}(\bar{z})$ $=\{i\,|\,i=1,\dots,n_{h}\},$
$\alpha(\bar{z})$ $=\{i\,|\,G_{i}(\bar{z})=0,H_{i}(\bar{z})>0\},$
$\gamma(\bar{z})$ $=\{i\,|\,G_{i}(\bar{z})>0,H_{i}(\bar{z})=0\},$
$\beta(\bar{z})$ $=\{i\,|\,G_{i}(\bar{z})=0,H_{i}(\bar{z})=0\}.$
A feasible point $\bar{z}$ is weakly stationary, if there exist multipliers $\bar{\lambda}=(\bar{\lambda}^{g},\bar{\lambda}^{h},\bar{\lambda}^{G},\bar{\lambda}^{H})$ with $\bar{\lambda}^{g}\geq 0$, such that
$0=\nabla f(\bar{z})+\sum_{i\in I_{g}(\bar{z})}\bar{\lambda}^{g}_{i}\nabla g_{i}(\bar{z})+\sum_{i\in I_{h}(\bar{z})}\bar{\lambda}_{i}^{h}\nabla h_{i}(\bar{z})\\ -\sum_{i\in\alpha(\bar{z})\cup\beta(\bar{z})}\bar{\lambda}^{G}_{i}\nabla G_{i}(\bar{z})-\sum_{i\in\gamma(\bar{z})\cup\beta(\bar{z})}\bar{\lambda}^{H}_{i}\nabla H_{i}(\bar{z}).$ (3)
Further, a weakly stationary point $\bar{z}$ is also
  • S-stationary (strongly stationary), if $\bar{\lambda}_{i}^{G},\bar{\lambda}_{i}^{H}\geq 0$ for all $i\in\beta(\bar{z})$;
  • M-stationary, if either $\bar{\lambda}_{i}^{G},\bar{\lambda}_{i}^{H}\geq 0$ or $\bar{\lambda}_{i}^{G}\bar{\lambda}_{i}^{H}=0$ for all $i\in\beta(\bar{z})$;
  • C-stationary, if $\bar{\lambda}_{i}^{G}\bar{\lambda}_{i}^{H}\geq 0$ for all $i\in\beta(\bar{z})$;
  • A-Stationary, if either $\bar{\lambda}_{i}^{G}\geq 0$ or $\bar{\lambda}_{i}^{H}\geq 0$ for all $i\in\beta(\bar{z})$.
1.1 First-Order Optimality Condition and Constraint Qualifications
A local minimizer $\bar{z}$ of MPCC (1) is a B-stationary point at which the following condition holds
$\nabla f(\bar{z})^{T}d\geq 0,\quad\forall d\in\mathcal{T}(\bar{z}),$ (4)
where $\mathcal{T}(\bar{z})$ is the tangent cone of the MPCC at the point $\bar{z}$. If the feasible region is regular at $\bar{z}$ in the sense of Clarke ([25, Definition 6.4][3, Section 1]), this condition is the same as
$\nabla f(\bar{z})\in\mathcal{T}(\bar{z})^{*},$ (5)
where $\mathcal{T}(\bar{z})^{*}$ is the dual cone of $\mathcal{T}(\bar{z})$. Without an analytic expression of the tangent cone, verifying these conditions directly is generally nontrivial. In practice, it is desirable to employ linearized cones to reconstruct the first-order condition (4) or (5). Constraint qualifications (CQs) play an important role in this task.
Standard linearization of $\mathcal{T}(\bar{z})$ can be carried out (see [7, Eqs. (10)-(11)]), by replacing the complementarity constraints $0\leq G(z)\perp H(z)\geq 0$ with
$G(z)\geq 0,\quad H(z)\geq 0,\quad G(z)^{T}H(z)=0.$
Linearizing these constraints gives that
$G_{i}(\bar{z})+\nabla G_{i}(\bar{z})^{T}d\geq 0,$ $i=1,\dots,m,$
$H_{i}(\bar{z})+\nabla H_{i}(\bar{z})^{T}d\geq 0,$ $i=1,\dots,m,$
$G_{i}(\bar{z})H_{i}(\bar{z})+H_{i}(\bar{z})\nabla G_{i}(\bar{z})^{T}d+G_{i}(\bar{z})\nabla H_{i}(\bar{z})^{T}d=0,$ $i=1,\dots,m,$
which, together with the index sets in (2), gives the following linearized tangent cone:
$\mathcal{T}^{\rm lin}(\bar{z})=\{d\,|$ $\nabla g_{i}(\bar{z})^{T}d\leq 0,$ $\forall i\in I_{g}(\bar{z}),$
$\nabla h_{i}(\bar{z})^{T}d=0,$ $\forall i\in I_{h}(\bar{z}),$
$\nabla G_{i}(\bar{z})^{T}d=0,$ $\forall i\in\alpha(\bar{z}),$
$\nabla H_{i}(\bar{z})^{T}d=0,$ $\forall i\in\gamma(\bar{z}),$
$\nabla G_{i}(\bar{z})^{T}d\geq 0,\;\nabla H_{i}(\bar{z})^{T}d\geq 0,$ $\forall i\in\beta(\bar{z})\}.$
Its dual cone is given by
$\mathcal{T}^{\rm lin}(\bar{z})^{*}$ $=\{w\,|\,w^{T}d\geq 0,\;\forall d\in\mathcal{T}^{\rm lin}(\bar{z})\}$
$\begin{aligned} \;=\{w\,|\,0=&\;w+\sum_{i\in I_{g}(\bar{z})}\bar{\lambda}_{i}^{g}\nabla g_{i}(\bar{z})+\sum_{i\in I_{h}(\bar{z})}\bar{\lambda}_{i}^{h}\nabla h_{i}(\bar{z})\\ ~&-\sum_{i\in\alpha(\bar{z})}\bar{\lambda}_{i}^{G}\nabla G_{i}(\bar{z})-\sum_{i\in\gamma(\bar{z})}\bar{\lambda}_{i}^{H}\nabla H_{i}(\bar{z})\\ ~&-\sum_{i\in\beta(\bar{z})}\bar{\lambda}_{i}^{G}\nabla G_{i}(\bar{z})-\sum_{i\in\beta(\bar{z})}\bar{\lambda}_{i}^{H}\nabla H_{i}(\bar{z});\\ ~&\bar{\lambda}_{i}^{g}\geq 0,\;\forall i\in I_{g}(\bar{z});\;\bar{\lambda}_{i}^{G}\geq 0,\bar{\lambda}_{i}^{H}\geq 0,\;\forall i\in\beta(\bar{z})\}.\end{aligned}$
The standard Abadie and Guignard constraint qualifications, NLP-ACQ and NLP-GCQ, assume $\mathcal{T}(\bar{z})=\mathcal{T}^{\rm lin}(\bar{z})$ and $\mathcal{T}(\bar{z})^{*}=\mathcal{T}^{\rm lin}(\bar{z})^{*}$, respectively, so that the conditions (4) and (5) can be rebuilt based on the linearized cone. This converts the first-order stationarity of MPCC (1) into that of the relaxed NLP
${\rm RNLP}:\quad\min$ $f(z)$ (6)
${\rm s.t.}$ $g(z)\leq 0,$
$h(z)=0,$
$G_{i}(z)=0,$ $i\in\alpha(\bar{z}),$
$H_{i}(z)=0,$ $i\in\gamma(\bar{z}),$
$G_{i}(z)\geq 0,\;H_{i}(z)\geq 0,$ $i\in\beta(\bar{z}),$
and thus justifies using the KKT conditions for RNLP, i.e., the S-stationarity condition, as a necessary first-order condition for the MPCC (see also [8, Theorem 4.1]). We should note that NLP-ACQ cannot hold if $\beta(\bar{z})\neq\emptyset$, because $\mathcal{T}(\bar{z})$ is a nonconvex cone in this case while $\mathcal{T}^{\rm lin}(\bar{z})$ is always a convex polyhedral cone.
Since NLP-CQs are usually too restrictive and not expected to hold in the presence of complementarity constraints, constraint qualifications tailored to MPCCs have been proposed. MPCC-ACQ [7] and MPCC-GCQ [9] extend the standard Abadie and Guignard constraint qualifications for NLPs to CQs that are tailored to the geometry of MPCCs. Instead of relating the tangent cone $\mathcal{T}(\bar{z})$ with $\mathcal{T}^{\rm lin}(\bar{z})$, they relate $\mathcal{T}(\bar{z})$ with the MPCC-linearized tangent cone $\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})$, which is given by:
$\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})=\{d\,|$ $\nabla g_{i}(\bar{z})^{T}d\leq 0,$ $\forall i\in I_{g}(\bar{z}),$
$\nabla h_{i}(\bar{z})^{T}d=0,$ $\forall i\in I_{h}(\bar{z}),$
$\nabla G_{i}(\bar{z})^{T}d=0,$ $\forall i\in\alpha(\bar{z}),$
$\nabla H_{i}(\bar{z})^{T}d=0,$ $\forall i\in\gamma(\bar{z}),$
$\nabla G_{i}(\bar{z})^{T}d\geq 0,$ $\forall i\in\beta(\bar{z}),$
$\nabla H_{i}(\bar{z})^{T}d\geq 0,$ $\forall i\in\beta(\bar{z}),$
$(\nabla G_{i}(\bar{z})^{T}d)\cdot(\nabla H_{i}(\bar{z})^{T}d)=0,$ $\forall i\in\beta(\bar{z})\}.$
MPCC-ACQ assumes $\mathcal{T}(\bar{z})=\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})$, then the condition (4) can be expressed as:
$\nabla f(\bar{z})^{T}d\geq 0,\quad\forall d\in\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z}).$
MPCC-GCQ assumes $\mathcal{T}(\bar{z})^{*}=\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})^{*}$, where
$\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})^{*}=\{w\,|w^{T}d\geq 0,\;\forall d\in\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})\},$
then the condition (5), if applicable, can be expressed as $\nabla f(\bar{z})\in\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})^{*}$. MPCC-GCQ is implied by MPCC-ACQ, but the converse is in general not true. Their relations are analogous to the relations between NLP-GCQ and NLP-ACQ. Examples showing that NLP-GCQ and MPCC-GCQ have a better chance to be satisfied, even if NLP-ACQ and MPCC-ACQ do not hold, can be found in [28, Example 1.3] and [9, Example 2.1]. Note that despite the fact that a tangent cone is not necessarily equal to the closure of its convex hull, their dual cones are the same. This offers the opportunity for NLP-GCQ and MPCC-GCQ to hold more generally. It has been established that under MPCC-GCQ, M-stationarity is a necessary first-order condition [9, Theorem 3.1]. In addition, Fritz John type M-stationarity has been derived at a local minimizer of an MPCC without requiring a constraint qualification [20, Theorem 3.1].
1.2 Degeneracy
To seek a solution of MPCC (1), various NLP-based schemes have been proposed. The original intention is to avoid dealing with the complementarity structure explicitly. In general, these schemes are designed to solve a sequence of regularized NLPs, yielding a sequence of stationary points $\{z^{k}\}$ that is hoped to approximate a solution of the MPCC. An important ingredient is to characterize conditions under which, as the regularization parameter vanishes or stabilizes, a limit point of a sequence $\{z^{k}\}$ is a stationary point of the MPCC in some sense. For some representative work, see [27, 11, 24, 23, 18, 19, 29, 10, 1].
A difficulty arises in establishing stationarity of a limit point that is degenerate (on the lower level), namely, a sequence $\{z^{k}\}\to\bar{z}$ at which $\beta(\bar{z})\neq\emptyset$. Fukushima and Pang [11] studied the behavior of a sequence $\{z^{k}\}$ that is composed of KKT points of NLPs formulated by smoothing the MPCC with perturbed Fischer-Burmeister functions. The condition of asymptotic weak nondegeneracy was proposed, meaning that for every $i\in\beta(\bar{z})$, $G_{i}(z^{k})$ and $H_{i}(z^{k})$ approach zero at the same order of magnitude. Under this condition and second-order necessary conditions at every $z^{k}$, together with MPCC linear independence constraint qualification (MPCC-LICQ) at $\bar{z}$, it has been proved that $\bar{z}$ is a B-stationary point of the MPCC [11, Theorem 3.1]. However, the condition of asymptotic weak nondegeneracy is hard to enforce in practice. Replacing this condition with upper level strict complementarity (ULSC), namely, $\bar{\lambda}_{i}^{G}\bar{\lambda}_{i}^{H}\neq 0$ for all $i\in\beta(\bar{z})$, Scholtes recovered B-stationarity of a limit point of a regularization scheme [27, Corollary 3.4]. Kadrani et al. developed a regularization method whose limit points were shown to be M-stationary under MPCC-LICQ, and S-stationary under additional assumption of asymptotic weak nondegeneracy [18]. The result on M-stationarity was later proved valid under weaker MPCC constant positive linear dependence (MPCC-CPLD) assumption [16]. Results under weaker assumptions also include, for example, that C-stationarity convergence of the method by Steffensen and Ulbrich under MPCC constant rank constraint qualification (MPCC-CRCQ) [29] and under MPCC-CPLD [15], and M-stationarity convergence of the method by Kanzow and Schwartz under MPCC-CPLD [19]. Theoretical and numerical comparison of some of these methods can be found in [16].
Besides diverse methods for reformulating complementarity constraints, many popular algorithmic frameworks in nonlinear programming have been exploited to deal with complementarity as well as the potential degeneracy. The sequential quadratic programming (SQP) methods applied to MPCCs were investigated in [10]. By introducing slack variables into the reformulation of general complementarity constraints, superlinear convergence to a S-stationary point was established under MPCC-LICQ and regularity conditions (Theorems 5.7 and 5.14 therein). An alternative SQP method that retained the superlinear convergence while relaxing some of the assumptions was analyzed in [2], where an adaptive elastic mode was invoked to enforce either feasibility of the QP subproblems or complementarity at the iterates (Theorems 4.5 and 4.6 therein). Interior-penalty methods for MPCCs were studied in [23]; global convergence to a S-stationary point was proved under MPCC-LICQ and a condition on the behavior of the penalty parameters (Theorem 3.4 and Corollary 3.5 therein); superlinear convergence to a S-stationary point was proved under certain regularity conditions (Theorem 4.5 therein); in particular, relations between interior-penalty and interior-relaxation methods were established, which allows to extend some convergence results derived for one approach to the other. Convergence of augmented Lagrangian methods was investigated under MPCC-LICQ [17, Theorem 3.2], where a limit point was proved to be S-stationary in the case of bounded multiplier sequence, and C-stationary in the presence of unbounded multiplier sequences. The results were improved in [1] for a second-order method (Theorem 3.2 therein), where S-stationarity was established under a weaker MPCC-relaxed constant positive linear dependence (MPCC-RCPLD) condition, and convergence in the presence of unbounded multipliers was proved to be M-stationary under MPCC-LICQ. Comparison of some existing augmented Lagrangian methods for MPCCs can be found in [14]. Finally, a detailed survey on these MPCC topics can be found in Kim et al. [21].
1.3 Outline
With a focus on the theoretically interesting case where $\beta(\bar{z})\neq\emptyset$, this paper investigates stationarity conditions for MPCCs and analyzes convergence behavior of MPCC methods. In Section 2, we propose a condition of piecewise M-stationarity and prove its equivalence to B-stationarity under MPCC-ACQ. In Section 3, we analyze convergence properties of the NCP-based Bounding Algorithm we proposed in [31]. An inequality variant of this algorithm offers an alternative viewpoint to interpret the behavior when approaching a non-strongly stationary solution, which together with the concept of piecewise M-stationarity leads to a cost reduction when verifying B-stationarity of a limit point (even if this point is not S-stationary). In Section 4, we discuss some practical issues for MPCC methods in their convergence to non-strongly stationary solutions. This shows an advantage of the NCP-based complementarity reformulations, that is, the structure of the generalized gradients of the underlying NCP function can prevent unbounded multipliers near these solutions.
2 Piecewise M-Stationarity
Given a feasible point $\bar{z}$ of MPCC (1), denote the set of all partitions of $\beta(\bar{z})$ by $\mathcal{P}(\beta(\bar{z}))=\{(\beta_{1},\beta_{2})\,|\,\beta_{1}\cup\beta_{2}=\beta(\bar{z}),\beta_{1}\cap\beta_{2}=\emptyset\}$. The so-called piecewise NLP problem defined on every partition $(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))$ is given by
${\rm NLP}_{(\beta_{1},\beta_{2})}:\quad\min$ $f(z)$ (7)
${\rm s.t.}$ $g(z)\leq 0,$
$h(z)=0,$
$G_{i}(z)=0,$ $i\in\alpha(\bar{z}),$
$H_{i}(z)=0,$ $i\in\gamma(\bar{z}),$
$G_{i}(z)=0,\;H_{i}(z)\geq 0,$ $i\in\beta_{1},$
$G_{i}(z)\geq 0,\;H_{i}(z)=0,$ $i\in\beta_{2}.$
Now we characterize piecewise M-stationarity, as well as the standard M-stationarity for purpose of contrast, based on the following Lagrangian function:
$\mathcal{L}(z,\lambda)=$ $f(z)+\sum_{i=1}^{n_{g}}\lambda_{i}^{g}g_{i}(z)+\sum_{i=1}^{n_{h}}\lambda_{i}^{h}h_{i}(z)-\sum_{i\in\alpha(z)}\lambda_{i}^{G}G_{i}(z)-\sum_{i\in\gamma(z)}\lambda_{i}^{H}H_{i}(z)$
$-\sum_{i\in\beta(z)}\lambda_{i}^{G}G_{i}(z)-\sum_{i\in\beta(z)}\lambda_{i}^{H}H_{i}(z).$
Definition 2.1.
Given a feasible point $\bar{z}$ of MPCC (1), we say that
  1. (a)
    $\bar{z}$ is M-stationary if there exist multipliers $\bar{\lambda}$ such that
    $\nabla_{z}\mathcal{L}(\bar{z},\bar{\lambda})=0,$
    $\bar{\lambda}_{i}^{g}\geq 0,\;\bar{\lambda}_{i}^{g}g_{i}(\bar{z})=0,$ $\quad\forall i\in\{1,\dots,n_{g}\},$
    $\text{either }\bar{\lambda}_{i}^{G},\bar{\lambda}_{i}^{H}\geq 0\text{ or }\bar{\lambda}_{i}^{G}\bar{\lambda}_{i}^{H}=0,$ $\quad\forall i\in\beta(\bar{z});$
  2. (b)
[…]
    $\nabla_{z}\mathcal{L}(\bar{z},\bar{\lambda})=0,$
    $\bar{\lambda}_{i}^{g}\geq 0,\;\bar{\lambda}_{i}^{g}g_{i}(\bar{z})=0,$ $\quad\forall i\in\{1,\dots,n_{g}\},$
    $\bar{\lambda}_{i}^{H}\geq 0,\;\bar{\lambda}_{i}^{H}H_{i}(\bar{z})=0,$ $\quad\forall i\in\beta_{1},$
    $\bar{\lambda}_{i}^{G}\geq 0,\;\bar{\lambda}_{i}^{G}G_{i}(\bar{z})=0,$ $\quad\forall i\in\beta_{2},$
    $\text{either }\bar{\lambda}_{i}^{G},\bar{\lambda}_{i}^{H}\geq 0\text{ or }\bar{\lambda}_{i}^{G}\bar{\lambda}_{i}^{H}=0,$ $\quad\forall i\in\beta(\bar{z}).$
[…]
2.1 Equivalence to B-stationarity
To establish the equivalence between piecewise M-stationarity and B-stationarity, we first show that, under MPCC-ACQ, B-stationarity implies piecewise M-stationarity.
Theorem 2.2.
Let $\bar{z}$ be a B-stationary point of MPCC (1) at which MPCC-ACQ holds. Then $\bar{z}$ is piecewise M-stationary.
Proof.
Recall that, under MPCC-ACQ, the B-stationarity condition (4) can be expressed as:
$\nabla f(\bar{z})^{T}d\geq 0,\quad\forall d\in\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z}),$ (8)
which is equivalent to that $d=0$ is the global minimizer of the following linear program with complementarity constraints (LPCC):
$\min$ $\nabla f(\bar{z})^{T}d$ (9)
${\rm s.t.}$ $\nabla g_{I}(\bar{z})^{T}d\leq 0,$
$\nabla h(\bar{z})^{T}d=0,$
$\nabla G_{\alpha}(\bar{z})^{T}d=0,$
$\nabla H_{\gamma}(\bar{z})^{T}d=0,$
$0\leq\nabla G_{\beta}(\bar{z})^{T}d\perp\nabla H_{\beta}(\bar{z})^{T}d\geq 0.$
Here $g_{I}$ denotes the constraints $\{g_{i}\,|\,\forall i\in I_{g}(\bar{z})\}$, and, similarly, $G_{\alpha},H_{\gamma},G_{\beta}$, and $H_{\beta}$ denote the constraints related to the index sets $\alpha(\bar{z}),\gamma(\bar{z})$, and $\beta(\bar{z})$. This LPCC is a combination of classic linear programs each defined on a partition $(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))$ as follows:
${\rm LP}_{(\beta_{1},\beta_{2})}:\quad\min$ ${\rm obj}(d)=\nabla f(\bar{z})^{T}d$ (10)
${\rm s.t.}$ $\nabla g_{I}(\bar{z})^{T}d\leq 0,$
$\nabla h(\bar{z})^{T}d=0,$
$\nabla G_{\alpha}(\bar{z})^{T}d=0,$
$\nabla H_{\gamma}(\bar{z})^{T}d=0,$
$\nabla G_{\beta_{1}}(\bar{z})^{T}d=0,\quad\nabla H_{\beta_{1}}(\bar{z})^{T}d\geq 0,$
$\nabla G_{\beta_{2}}(\bar{z})^{T}d\geq 0,\quad\nabla H_{\beta_{2}}(\bar{z})^{T}d=0.$
[…]
$\nabla f(\bar{z})^{T}d\geq 0,\quad\forall d\in\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z}),$ (11)
where
$\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})=\{d\,|$ $\nabla g_{i}(\bar{z})^{T}d\leq 0,$ $\forall i\in I_{g}(\bar{z}),$
$\nabla h_{i}(\bar{z})^{T}d=0,$ $\forall i\in I_{h}(\bar{z}),$
$\nabla G_{i}(\bar{z})^{T}d=0,$ $\forall i\in\alpha(\bar{z}),$
$\nabla H_{i}(\bar{z})^{T}d=0,$ $\forall i\in\gamma(\bar{z}),$
$\nabla G_{i}(\bar{z})^{T}d=0,\;\nabla H_{i}(\bar{z})^{T}d\geq 0,$ $\forall i\in\beta_{1},$
$\nabla G_{i}(\bar{z})^{T}d\geq 0,\;\nabla H_{i}(\bar{z})^{T}d=0,$ $\forall i\in\beta_{2}\}.$
[…]
$\nabla f(\bar{z})\in\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})^{*},\quad\forall(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z})),$ (12)
where
$\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})^{*}=\{w\,|\,0=w+\nabla g_{I}(\bar{z})\eta_{I}^{g}+\nabla h(\bar{z})\eta^{h}-\nabla G_{\alpha}(\bar{z})\eta_{\alpha}^{G}-\nabla H_{\gamma}(\bar{z})\eta_{\gamma}^{H}\\ \begin{aligned} &-\nabla G_{\beta_{1}}(\bar{z})\eta_{\beta_{1}}^{G}-\nabla H_{\beta_{1}}(\bar{z})\eta_{\beta_{1}}^{H}-\nabla G_{\beta_{2}}(\bar{z})\eta_{\beta_{2}}^{G}-\nabla H_{\beta_{2}}(\bar{z})\eta_{\beta_{2}}^{H},\\ &\eta_{I}^{g},\eta_{\beta_{1}}^{H},\eta_{\beta_{2}}^{G}\geq 0\}.\end{aligned}$ (13)
On the other hand, since $\bar{z}$ satisfies the first-order necessary condition (8) for optimality, there exists a scalar $\lambda_{0}\geq 0$ and multipliers $\lambda_{I}^{g}\geq 0,\lambda^{h},\lambda_{\alpha}^{G},\lambda_{\gamma}^{H},\zeta$, satisfying the following Fritz John condition ([5, Theorem 6.1.1],[26, Lemma 1 and proof]):
$0$ $\in$ $\lambda_{0}\nabla f(\bar{z})+\nabla g_{I}(\bar{z})\lambda_{I}^{g}+\nabla h(\bar{z})\lambda^{h}-\nabla G_{\alpha}(\bar{z})\lambda_{\alpha}^{G}-\nabla H_{\gamma}(\bar{z})\lambda_{\gamma}^{H}$ (14)
$-\sum_{i\in\beta(\bar{z})}\zeta_{i}\,{\rm conv}\{\nabla G_{i}(\bar{z}),\nabla H_{i}(\bar{z})\},$
[…]
Let $\theta_{i}\nabla G_{i}(\bar{z})+(1-\theta_{i})\nabla H_{i}(\bar{z})$ with some $\theta_{i}\in[0,1]$ be an element of the convex hull, then we have (see also [28, Section 2.2])
$0$ $=$ $\lambda_{0}\nabla f(\bar{z})+\nabla g_{I}(\bar{z})\lambda_{I}^{g}+\nabla h(\bar{z})\lambda^{h}-\nabla G_{\alpha}(\bar{z})\lambda_{\alpha}^{G}-\nabla H_{\gamma}(\bar{z})\lambda_{\gamma}^{H}$ (15)
$-\sum_{i\in\beta_{1}}\underbrace{\zeta_{i}\theta_{i}}_{\lambda_{i}^{G}}\nabla G_{i}(\bar{z})-\sum_{i\in\beta_{1}}\underbrace{\zeta_{i}(1-\theta_{i})}_{\lambda_{i}^{H}}\nabla H_{i}(\bar{z})$
$-\sum_{i\in\beta_{2}}\underbrace{\zeta_{i}\theta_{i}}_{\lambda_{i}^{G}}\nabla G_{i}(\bar{z})-\sum_{i\in\beta_{2}}\underbrace{\zeta_{i}(1-\theta_{i})}_{\lambda_{i}^{H}}\nabla H_{i}(\bar{z}).$
For every $(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))$, this system has a solution with $\lambda_{0}=1$ and $\lambda_{I}^{g},\lambda_{\beta_{1}}^{H},\lambda_{\beta_{2}}^{G}\geq 0$, as indicated by the conditions (12)-(13). To be specific, because $\bar{z}$ satisfies the first-order condition (8), we have that, for every $(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))$, the relations (12)-(13) and (15) hold simultaneously. Letting the multipliers in (12)-(13) take the form as in (15), we obtain that (15) has a solution with $\lambda_{0}=1$ (without loss of generality) and
$\begin{gathered}\lambda_{I}^{g}=\eta_{I}^{g}\geq 0,\quad\lambda^{h}=\eta^{h},\quad\lambda_{\alpha}^{G}=\eta_{\alpha}^{G},\quad\lambda_{\gamma}^{H}=\eta_{\gamma}^{H},\\ \begin{aligned} &\lambda_{i}^{G}=\zeta_{i}\theta_{i}=\eta_{i}^{G},&&\lambda_{i}^{H}=\zeta_{i}(1-\theta_{i})=\eta_{i}^{H}\geq 0,\quad&&\forall i\in\beta_{1},\\ &\lambda_{i}^{G}=\zeta_{i}\theta_{i}=\eta_{i}^{G}\geq 0,&&\lambda_{i}^{H}=\zeta_{i}(1-\theta_{i})=\eta_{i}^{H},\quad&&\forall i\in\beta_{2}.\end{aligned}\end{gathered}$
It follows from $\lambda_{\beta_{1}}^{H},\lambda_{\beta_{2}}^{G}\geq 0$ that
$i$ $\in\beta_{1}\left\{\begin{aligned} \zeta_{i}\geq 0\implies&\exists\theta_{i}\in[0,1],&&\lambda_{i}^{G}\geq 0,\;\lambda_{i}^{H}\geq 0,\\ \zeta_{i}<0\implies&\theta_{i}=1,&&\lambda_{i}^{G}=\zeta_{i}<0,\;\lambda_{i}^{H}=0,\end{aligned}\right$. (16)
$i$ $\in\beta_{2}\left\{\begin{aligned} \zeta_{i}\geq 0\implies&\exists\theta_{i}\in[0,1],&&\lambda_{i}^{G}\geq 0,\;\lambda_{i}^{H}\geq 0,\\ \zeta_{i}<0\implies&\theta_{i}=0,&&\lambda_{i}^{G}=0,\;\lambda_{i}^{H}=\zeta_{i}<0.\end{aligned}\right$.
This also shows that, for every $(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))$, either $\lambda_{i}^{G},\lambda_{i}^{H}\geq 0$ or $\lambda_{i}^{G}\lambda_{i}^{H}=0$ for all $i\in\beta(\bar{z})$. This completes the proof.
The following theorem ensures that piecewise M-stationarity implies B-stationarity.
Theorem 2.3.
Let $\bar{z}$ be a feasible point of MPCC (1). If $\bar{z}$ is piecewise M-stationary, then $\bar{z}$ is B-stationary.
Proof.
[…]
Corollary 2.4.
Let $\bar{z}$ be a feasible point of MPCC (1) at which MPCC-ACQ holds. Then $\bar{z}$ is B-stationary if and only if $\bar{z}$ is piecewise M-stationary.
Proof.
The necessary part and sufficient part follow from Theorem 2.2 and Theorem 2.3, respectively. ∎
The MPCC-ACQ assumption in Corollary 2.4 may not be relaxed. To see this, we first note that piecewise M-stationarity of $\bar{z}$ needs the existence of KKT multipliers at $\bar{z}$ for each of the piecewise NLPs. Then, according to [13, Lemma 4.2], we obtain that the condition (12) holds. It follows that $\nabla f(\bar{z})^{T}d\geq 0$ for all $d\in\bigcup_{(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))}\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})=\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})$, whose equivalence to B-stationarity requires the MPCC-ACQ assumption. Furthermore, noting that the sufficient part of this corollary does not rely on MPCC-ACQ, we therefore know that the MPCC-ACQ assumption for the necessary part, i.e., Theorem 2.2, may not be relaxed.
2.2 Example: Piecewise M-Stationarity Illustration
Problem scholtes4 from the MacMPEC collection [22] is given by
[…]
${\rm s.t.}$ $-4z_{1}+z_{3}\leq 0,$ $\lambda_{1}$
$-4z_{2}+z_{3}\leq 0,$ $\lambda_{2}$
$0\leq z_{1}\perp z_{2}\geq 0.$ $\sigma_{1},\sigma_{2}$
The global minimizer $\bar{z}=(z_{1},z_{2},z_{3})=(0,0,0)$ is B-stationarity, at which MPCC-MFCQ holds and $\beta(\bar{z})=\{1\}$. The multipliers at $\bar{z}$ satisfy
$\lambda_{1}+\lambda_{2}$ $=1,$
$\sigma_{1}+\sigma_{2}$ $=-2,$
[…]
Moreover, this example shows that the conditions (12)-(13) alone are not sufficient to relate $\nabla f(\bar{z})$ to the MPCC. We have from (12)-(13) that
$\nabla f(\bar{z})=\begin{bmatrix}1\\ 1\\ -1\end{bmatrix}$ $\in\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})^{*}=\left\{w|\begin{bmatrix}w_{1}\\ w_{2}\\ w_{3}\end{bmatrix}=\begin{bmatrix}4\lambda_{1}+\sigma_{1}\\ 4\lambda_{2}+\sigma_{2}\\ -\lambda_{1}-\lambda_{2}\end{bmatrix},\lambda_{1},\lambda_{2},\sigma_{2}\geq 0\right\},$ (17a)
$\nabla f(\bar{z})=\begin{bmatrix}1\\ 1\\ -1\end{bmatrix}$ $\in\mathcal{T}_{(\hat{\beta}_{1},\hat{\beta}_{2})}^{\rm lin}(\bar{z})^{*}=\left\{w|\begin{bmatrix}w_{1}\\ w_{2}\\ w_{3}\end{bmatrix}=\begin{bmatrix}4\lambda_{1}+\sigma_{1}\\ 4\lambda_{2}+\sigma_{2}\\ -\lambda_{1}-\lambda_{2}\end{bmatrix},\lambda_{1},\lambda_{2},\sigma_{1}\geq 0\right\},$ (17b)
which relates $\nabla f(\bar{z})$ to the piecewise NLPs. However, we still cannot relate $\nabla f(\bar{z})$ to the MPCC because $\nabla f(\bar{z})\notin\mathcal{T}_{\rm MPCC}^{\rm lin}(\bar{z})^{*}$, namely,
$\nabla f(\bar{z})=\begin{bmatrix}1\\ 1\\ -1\end{bmatrix}$ $\notin\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})^{*}\cap\mathcal{T}_{(\hat{\beta}_{1},\hat{\beta}_{2})}^{\rm lin}(\bar{z})^{*}$
$=\left\{w|\begin{bmatrix}w_{1}\\ w_{2}\\ w_{3}\end{bmatrix}=\begin{bmatrix}4\lambda_{1}+\sigma_{1}\\ 4\lambda_{2}+\sigma_{2}\\ -\lambda_{1}-\lambda_{2}\end{bmatrix},\lambda_{1},\lambda_{2},\sigma_{1},\sigma_{2}\geq 0\right\}.$
The right side indicates that $\nabla f(\bar{z})$ does not locate in the intersection except when $\bar{z}$ is S-stationary; otherwise this fails to relate $\nabla f(\bar{z})$ to the MPCC, and we cannot say more about the multipliers at $\bar{z}$ beyond their separate existence on each piece.
In fact, the condition (15) gives a further characterization on the stationarity at $\bar{z}$. That is, for every $i\in\beta(\bar{z})$, the multipliers for $\nabla G_{i}(\bar{z})$ and $\nabla H_{i}(\bar{z})$ are in the restricted form of $\zeta_{i}\theta_{i}$ and $\zeta_{i}(1-\theta_{i})$ with $\theta_{i}\in[0,1]$, respectively. Letting the multipliers $\sigma_{1},\sigma_{2}$ in (17) take this form, we obtain that
$\nabla f(\bar{z})=\begin{bmatrix}1\\ 1\\ -1\end{bmatrix}$ $\in\left\{w|\begin{bmatrix}w_{1}\\ w_{2}\\ w_{3}\end{bmatrix}=\begin{bmatrix}4\lambda_{1}+\zeta\theta\\ 4\lambda_{2}+\zeta(1-\theta)\\ -\lambda_{1}-\lambda_{2}\end{bmatrix},\lambda_{1},\lambda_{2},\zeta(1-\theta)\geq 0\right\},$ (18a)
$\nabla f(\bar{z})=\begin{bmatrix}1\\ 1\\ -1\end{bmatrix}$ $\in\left\{w|\begin{bmatrix}w_{1}\\ w_{2}\\ w_{3}\end{bmatrix}=\begin{bmatrix}4\lambda_{1}+\zeta\theta\\ 4\lambda_{2}+\zeta(1-\theta)\\ -\lambda_{1}-\lambda_{2}\end{bmatrix},\lambda_{1},\lambda_{2},\zeta\theta\geq 0\right\}.$ (18b)
Comparing (17) and (18), we note that the cones in the latter restrict the choices of $\sigma_{1}$ and $\sigma_{2}$, and therefore prevent (17a) and (17b) holding simultaneously in the case of $\zeta<0$. Choosing $\theta$ according to (16) reveals complementarity of the multipliers $\sigma_{1}$ and $\sigma_{2}$, and therefore piecewise M-stationarity of $\bar{z}$ for the MPCC.
2.3 Example: Piecewise M-Stationarity Failure at an M-stationary Point
This example shows that piecewise M-stationarity is a stronger condition than M-stationarity. Consider the problem given by
[…]
${\rm s.t.}$ $0\leq z_{1}\perp z_{2}\geq 0.$ $\sigma_{1},\sigma_{2}$
MPCC-LICQ holds at every feasible point of this problem. The global minimizer is $z^{*}=(1,0)$, which is S-stationary, and thus B-stationary and trivially piecewise M-stationary since $\beta(\bar{z})=\emptyset$.
[…]
$\nabla f(\bar{z})=\begin{bmatrix}-2\\ 0\end{bmatrix}\notin\mathcal{T}_{(\hat{\beta}_{1},\hat{\beta}_{2})}^{\rm lin}(\bar{z})^{*}=\{w\,|\,w_{1}=\sigma_{1},w_{2}=\sigma_{2},\sigma_{1}\geq 0\},$
which indicates that $\bar{z}$ is not piecewise M-stationary.
2.4 Example: Piecewise M-Stationarity Failure at a Fritz John Point
Consider the following problem:
[…]
${\rm s.t.}$ $z_{2}^{2}\leq 0,$ $\lambda$
$0\geq z_{1}\perp z_{2}\leq 0,$ $\sigma_{1},\sigma_{2}$
which searches for the minimal distance between points $(z_{1},z_{2})$ and (1,1)(1,-1), along the negative axis of $z_{1}$. The solution is $\bar{z}=(0,0)$, where the multipliers are $(\sigma_{1},\sigma_{2})=(2,-2)$, and therefore $\bar{z}$ is A-stationary.
In this example, MPCC-GCQ, and therefore MPCC-ACQ, fails at $\bar{z}$. For partitions $(\beta_{1},\beta_{2})=(\{1\},\emptyset)$ and $(\hat{\beta}_{1},\hat{\beta}_{2})=(\emptyset,\{1\})$, we have that
$\nabla f(\bar{z})=\begin{bmatrix}-2\\ 2\end{bmatrix}$ $\notin\mathcal{T}_{(\beta_{1},\beta_{2})}^{\rm lin}(\bar{z})^{*}=\{w|w_{1}=-\sigma_{1},w_{2}=-\sigma_{2},\lambda\geq 0,\sigma_{2}\geq 0\},$
$\nabla f(\bar{z})=\begin{bmatrix}-2\\ 2\end{bmatrix}$ $\in\mathcal{T}_{(\hat{\beta}_{1},\hat{\beta}_{2})}^{\rm lin}(\bar{z})^{*}=\{w|w_{1}=-\sigma_{1},w_{2}=-\sigma_{2},\lambda\geq 0,\sigma_{1}\geq 0\},$
[…]
$0=\lambda_{0}\begin{bmatrix}-2\\ 2\end{bmatrix}+\lambda\begin{bmatrix}0\\ 0\end{bmatrix}+\zeta\theta\begin{bmatrix}1\\ 0\end{bmatrix}+\zeta(1-\theta)\begin{bmatrix}0\\ 1\end{bmatrix}$,
which implies that $\lambda_{0}=\zeta=0$ and $\lambda>0$ can be chosen arbitrarily.
2.5 Distinction from Extended M-stationarity
So far we have shown that, under a mild assumption, piecewise M-stationarity is an equivalent concept to B-stationarity and therefore a stronger concept than the standard M-stationarity. A concept that has strong similarities to piecewise M-stationarity is extended M-stationarity, which was proposed by Gfrerer [12] and also discussed in Kim et al. [21]. Extended M-stationarity and its equivalence to B-stationarity were investigated in the background of disjunctive programs (problems in a more general form than MPCCs), based on the concepts of generalized differentiation (such as metric regularity, subregularity, mixed regularity/subregularity, and their directional versions). We also note the following subtle differences between extended and piecewise M-stationarity concepts, used as equivalence to B-stationarity.
  • Since extended M-stationarity is usually difficult to verify in practice [12, Page 924], Definition 4.2 in [12] introduces the concept of strong M-stationarity as a bridge between M-stationarity and S-stationarity. Under an assumption of linearly independent gradients of some MPCC working set of constraints, [12, Theorem 4.3] shows that an extended M-stationary point is strongly M-stationary.
  • A summary of extended M-stationarity and its relation to other properties is given on [12, Page 928] and in [21, Figure 12.2]. Here an extended M-stationary point is also B-stationary, and a B-stationary point that satisfies GGCQ (which we call MPCC-GCQ) is also extended M-stationary.
On the other hand, piecewise M-stationarity requires the stronger assumption of MPCC-ACQ to establish the equivalence to B-stationarity, and it is straightforward to verify through M-stationarity of each branch of the LPCC at $\bar{z}$. It is not clear that these different conditions can be reconciled easily to the ones in [12]. Therefore, these provide a distinction between extended and piecewise M-stationarity.
3 Convergence of NCP-based Bounding Methods
The results of Section 2 are independent of algorithms designed for solving MPCCs. In the sequel, we investigate convergence properties of the NCP-based bounding methods we proposed in [31], without assuming MPCC-LICQ.
3.1 Brief Review of a Bounding Scheme
In [31], we proposed an algorithm to seek a solution of MPCC (1) by solving a sequence of NLP problems of the form
[…]
${\rm s.t.}$ $g(z)\leq 0,$ $u^{g}$
$h(z)=0,$ $u^{h}$
$\Phi^{\epsilon}_{i}(z)+p_{i}=0,\;i=1,\dots,m,$ $u^{\Phi}_{i}$
where
$\Phi^{\epsilon}_{i}(z)=\frac{1}{2}\left(G_{i}(z)+H_{i}(z)-\sqrt{(G_{i}(z)-H_{i}(z))^{2}+\epsilon^{2}}\right)$ (20)
is a smoothed NCP function (due to [4]), the smoothing factor $\epsilon>0$, and the parameter $p_{i}$ is adjusted adaptively (to take a value of zero or $\epsilon/2$). We define the Lagrangian for the problem BA($\epsilon$) as
$\mathcal{L}(z,u)=f(z)+\sum_{i\in I_{g}(z)}u_{i}^{g}g_{i}(z)+\sum_{i\in I_{h}(z)}u_{i}^{h}h_{i}(z)-\sum_{i=1}^{m}u_{i}^{\Phi}(\Phi_{i}^{\epsilon}(z)+p_{i}).$ (21)
As $\epsilon\to 0$, a sequence of KKT points of BA($\epsilon$) tends to a limit point. Main results of this method are summarized below.
  • Feasibility: The smoothed NCP function (20) is used to approximate the complementarity constraints in MPCC (1), and the largest difference between them is $\epsilon/2$ [31, Proposition 1.7]. When $\epsilon>0$, every feasible point $z$ of BA($\epsilon$) satisfies
    $\Phi^{\epsilon}_{i}(z)+p_{i}=0\quad\Leftrightarrow$ (22)
    $G_{i}(z)+p_{i}>0,\;H_{i}(z)+p_{i}>0,\;(G_{i}(z)+p_{i})(H_{i}(z)+p_{i})=\left(\tfrac{\epsilon}{2}\right)^{2},$
    whose limit at $\epsilon=0$ (thus $p_{i}=0$) recovers $0\leq G_{i}(z)\perp H_{i}(z)\geq 0$.
  • Sensitivity: At a KKT point $z(p)$ of BA($\epsilon$), the sensitivities $\tfrac{{\rm d}f(z(p))}{{\rm d}p_{i}}$ are given by $u^{\Phi}_{i}$ for $i=1,\dots,m$, provided that NLP-LICQ, second-order sufficient conditions, and strict complementarity hold at $z(p)$. This observation throws some light on the design of the Bounding Algorithm, which takes advantage of the sensitivities to yield an efficient isolation of a solution to the MPCC.
  • Convergence: The following convergence results have been established for the Bounding Algorithm applied to BA($\epsilon$).
    1. (i)
      Suppose that MPCC-LICQ holds at a feasible point of the MPCC. Then in a neighborhood of this point, NLP-LICQ holds at every feasible point of BA($\epsilon$), whenever $\epsilon>0$ is sufficiently small [31, Theorem 3.1].
    2. (ii)
      Suppose that a sequence of KKT points of problems BA($\epsilon$) tends to a limit point as $\epsilon\to 0$, at which MPCC-LICQ holds. Then the limit point is C-stationary [31, Theorem 3.3].
    3. (iii)
      In addition, suppose that the reduced Hessian of the Lagrangian at each of the KKT points of problems BA($\epsilon$) is bounded below when $\epsilon>0$ is sufficiently small. Then the limit point is M-stationary [31, Theorem 3.5].
3.2 Bounding Algorithm
The main idea of the Bounding Algorithm is given below to facilitate the later analysis.
Since $\epsilon/2\leq\Phi_{i}^{\epsilon}(z)\leq 0$ encloses the complementarity condition $0\leq G_{i}(z)\perp H_{i}(z)\geq 0$ [31, Proposition 1.7], we can approximate the complementarity with the equation $\Phi_{i}^{\epsilon}(z)+p_{i}=0$ by letting $p_{i}\in[0,\epsilon/2]$ and $\epsilon\to 0$. For any parameters $p_{i},p_{i}^{\prime}\in[0,\epsilon/2]$ with $\epsilon>0,i=1,\dots,m$, and the corresponding solutions $z(p)$ and $z(p^{\prime})$ to BA($\epsilon$), it is straightforward to show that
$f(z(p^{\prime}))=f(z(p))+\left[\frac{{\rm d}f(z(p))}{{\rm d}p}\right]^{T}(p^{\prime}-p)+O(\|p^{\prime}-p\|^{2}).$
Noting that the sensitivities $\tfrac{{\rm d}f(z(p))}{{\rm d}p}$ are given by $u^{\Phi}$, we have that
$f(z(p))-\frac{\epsilon}{2}\sum_{i=1}^{m}|u_{i}^{\Phi}(p)|-|O(\epsilon^{2})|\leq f(z(p^{\prime}))\leq f(z(p))+\frac{\epsilon}{2}\sum_{i=1}^{m}|u_{i}^{\Phi}(p)|+|O(\epsilon^{2})|.$
This relation explains the estimates of the objective by the Bounding Algorithm. In order to seek a solution to the MPCC, the problem BA($\epsilon^{k}$) is solved to a local solution $z^{k}$, for a sequence of smoothing parameters $\{\epsilon^{k}\}\to 0$. At each solution $z^{k}=z(p^{k})$, we take advantage of the sensitivities $u^{\Phi,k}=u^{\Phi}(p^{k})$ to adjust the parameters $p^{k}$, with the aim of improving the objective in the subsequent solution. As shown in Step 3 of the algorithm, in the case of $u_{i}^{\Phi,k}>0$ (or $u_{i}^{\Phi,k}<0$), increasing (or decreasing) $p_{i}^{k}$ may lead to decrease in the objective function, and the corresponding indices make up a set $P_{0}$ (or $P_{\epsilon}$). Step 4 then adjusts the parameters for the next solution; $p_{i}^{k+1}$ takes the extreme values for $i\in P_{0}\cup P_{\epsilon}$; otherwise, $p_{i}^{k+1}$ is set to $\kappa p_{i}^{k}$ for $i\notin P_{0}\cup P_{\epsilon}$, where the constant $\kappa\in(0,1)$. When $\epsilon^{k}>0$ is sufficiently small, $z(p^{k})$ is an $\epsilon^{k}$-approximate solution to the MPCC, which includes an $O((\epsilon^{k})^{2})$ correction arising from the parameters adjustment [31, Proposition 2.1].
Algorithm A Bounding Algorithm for MPCCs
Initialization: Specify initial smoothing factor $\epsilon^{0}>0$, reducing factor $\kappa\in(0,1)$, initial point $z^{0}$, solution tolerance $\epsilon_{\rm tol}>0$. Set initial parameters $p^{0}\leftarrow 0$, counter $k\leftarrow 0$.
Main loop: While $\epsilon^{k}\geq\epsilon_{\rm tol}$, do the following.
  1. 1.
    Solve the problem BA($\epsilon^{k}$) with parameters $p^{k}$, to obtain a stationary point $z^{k}$ and multipliers $u^{k}=(u^{g,k},u^{h,k},u^{\Phi,k})$.
  2. 2.
    Approximate the upper bound of the MPCC’s objective value with
    $f^{\rm up}=f(z^{k})+\epsilon^{k}\sum_{i=1}^{m}|u^{\Phi,k}_{i}|.$
  3. 3.
    Approximate the lower bound of the MPCC’s objective value as follows. Define the index sets
    $P_{0}$ $=\{i\,|\,p_{i}^{k}=0\text{ and }u^{\Phi,k}_{i}>0\},$
    $P_{\epsilon}$ $=\{i\,|\,p_{i}^{k}=\epsilon^{k}/2\text{ and }u^{\Phi,k}_{i}<0\}.$
    Then the following settings would reduce $f(z^{k})$:
    $p_{i}^{k}$ $\leftarrow\epsilon^{k}/2,$ $\forall i\in P_{0},$
    $p_{i}^{k}$ $\leftarrow 0,$ $\forall i\in P_{\epsilon}.$
    The objective with the adjustment of $p^{k}$ would approximately be
    $f^{\rm low}=f(z^{k})-\epsilon^{k}\sum_{i\in P_{0}\cup P_{\epsilon}}|u^{\Phi,k}_{i}|.$
  4. 4.
    Update the parameters $\epsilon$ and $p$. Set $\epsilon^{k+1}\leftarrow\kappa\epsilon^{k}$, and
    $p_{i}^{k+1}=\left\{\begin{aligned} &\epsilon^{k+1}/2,&&i\in P_{0},\\ &0,&&i\in P_{\epsilon},\\ &\kappa p_{i}^{k},&&{\rm otherwise}.\end{aligned}\right$.
  5. 5.
    Set $k\leftarrow k+1$ and go to Step 1.
3.3 Derivatives of $\Phi^{\epsilon}$
Derivatives of the function $\Phi^{\epsilon}$ defined in (20) are derived in [31]. The following gives the main results for their later use.
With $\epsilon>0$, the first and second derivatives of the function $\Phi_{i}^{\epsilon}(z)$ in (20) are given by
$\nabla_{G}\Phi^{\epsilon}_{i}(z)$ $=\frac{1}{2}-\frac{G_{i}(z)-H_{i}(z)}{2\sqrt{(G_{i}(z)-H_{i}(z))^{2}+\epsilon^{2}}},$
$\nabla_{H}\Phi^{\epsilon}_{i}(z)$ $=\frac{1}{2}+\frac{G_{i}(z)-H_{i}(z)}{2\sqrt{(G_{i}(z)-H_{i}(z))^{2}+\epsilon^{2}}},$
$\nabla_{GG}\Phi^{\epsilon}_{i}(z)=\nabla_{HH}\Phi^{\epsilon}_{i}(z)$ $=\frac{-\epsilon^{2}}{2[(G_{i}(z)-H_{i}(z))^{2}+\epsilon^{2}]^{3/2}},$
$\nabla_{GH}\Phi^{\epsilon}_{i}(z)=\nabla_{HG}\Phi^{\epsilon}_{i}(z)$ $=\frac{\epsilon^{2}}{2[(G_{i}(z)-H_{i}(z))^{2}+\epsilon^{2}]^{3/2}}.$
Let $z$ satisfy $\Phi_{i}^{\epsilon}(z)+p_{i}=0$ with $\epsilon>0$. It follows from (22) that
$\sqrt{(G_{i}(z)-H_{i}(z))^{2}+\epsilon^{2}}=\sqrt{((G_{i}(z)+p_{i})-(H_{i}(z)+p_{i}))^{2}+\epsilon^{2}}\\ \begin{aligned} &=\sqrt{(G_{i}(z)+p_{i})^{2}+(H_{i}(z)+p_{i})^{2}+2(G_{i}(z)+p_{i})(H_{i}(z)+p_{i})}\\ &=|G_{i}(z)+H_{i}(z)+2p_{i}|=G_{i}(z)+H_{i}(z)+2p_{i}.\end{aligned}$
Using this and $(G_{i}(z)+p_{i})(H_{i}(z)+p_{i})=(\epsilon/2)^{2}$, we can rephrase the above derivatives as
$\nabla_{G}\Phi^{\epsilon}_{i}(z)$ $=\frac{H_{i}(z)+p_{i}}{G_{i}(z)+H_{i}(z)+2p_{i}},$ (23)
$\nabla_{H}\Phi^{\epsilon}_{i}(z)$ $=\frac{G_{i}(z)+p_{i}}{G_{i}(z)+H_{i}(z)+2p_{i}},$
$\nabla_{GG}\Phi^{\epsilon}_{i}(z)=\nabla_{HH}\Phi^{\epsilon}_{i}(z)$ $=\frac{-2(G_{i}(z)+p_{i})(H_{i}(z)+p_{i})}{(G_{i}(z)+H_{i}(z)+2p_{i})^{3}},$
$\nabla_{GH}\Phi^{\epsilon}_{i}(z)=\nabla_{HG}\Phi^{\epsilon}_{i}(z)$ $=\frac{2(G_{i}(z)+p_{i})(H_{i}(z)+p_{i})}{(G_{i}(z)+H_{i}(z)+2p_{i})^{3}}.$
3.4 C-Stationarity Convergence
Let a sequence $\{z^{k}\}\to\bar{z}$ as $\epsilon^{k}\to 0$, where every $z^{k}$ is a KKT point of BA($\epsilon^{k}$), namely, there exist multipliers $u^{k}=(u^{g,k},u^{h,k},u^{\Phi,k})$ with $u^{g,k}\geq 0$ such that
$0=\nabla f(z^{k})+\sum_{i\in I_{g}(z^{k})}u_{i}^{g,k}\nabla g_{i}(z^{k})+\sum_{i\in I_{h}(z^{k})}u_{i}^{h,k}\nabla h_{i}(z^{k})-\sum_{i=1}^{m}u^{\Phi,k}_{i}\nabla\Phi_{i}^{\epsilon}(z^{k}).$ (24)
The following establishes C-stationarity of $\bar{z}$.
Theorem 3.1.
For a sequence of positive scalars $\epsilon^{k}\to 0$, apply the Bounding Algorithm to BA($\epsilon^{k}$). Assume this generates a sequence $\{z^{k}\}\to\bar{z}$, where every $z^{k}$ is a KKT point of BA($\epsilon^{k}$) and suppose MPCC-MFCQ holds at $\bar{z}$. Then the following statements hold.
  1. 1.
    For every $\epsilon^{k}>0$ sufficiently small, the NLP multipliers $u^{k}$ are bounded.
  2. 2.
    The point $\bar{z}$ is a C-stationary point of MPCC (1) with multipliers which satisfy
    $\bar{\lambda}^{g}$ $=\bar{u}^{g}=\lim_{k\to\infty}u^{g,k},$ (25)
    $\bar{\lambda}^{h}$ $=\bar{u}^{h}=\lim_{k\to\infty}u^{h,k},$
    $\bar{\lambda}_{i}^{G}$ $=\left\{\begin{aligned} &\bar{u}^{\Phi}_{i}=\lim_{k\to\infty}u_{i}^{\Phi,k},&&i\in\alpha(\bar{z})\\ &\bar{u}^{\Phi}_{i}\theta_{i},&&i\in\beta(\bar{z}),\end{aligned}\right$.
    $\bar{\lambda}_{i}^{H}$ $=\left\{\begin{aligned} &\bar{u}^{\Phi}_{i}=\lim_{k\to\infty}u_{i}^{\Phi,k},&&i\in\gamma(\bar{z})\\ &\bar{u}^{\Phi}_{i}(1-\theta_{i}),&&i\in\beta(\bar{z}),\end{aligned}\right$.
    where $\theta_{i}\in[0,1]$.
  3. 3.
    The point $\bar{z}$ is S-stationary and therefore B-stationary if $\bar{u}_{i}^{\Phi}\geq 0$ for all $i\in\beta(\bar{z})$.
Proof.
Consider the stationarity condition (24) at a KKT point $z^{k}$ of BA($\epsilon^{k}$).
Derivatives in the limit. When $\epsilon^{k}>0$, the gradient of $\Phi_{i}^{\epsilon}$ is given by
$\nabla\Phi_{i}^{\epsilon}(z^{k})=$ $\nabla_{G}\Phi_{i}^{\epsilon}(z^{k})\nabla G_{i}(z^{k})+\nabla_{H}\Phi_{i}^{\epsilon}(z^{k})\nabla H_{i}(z^{k})$
$=$ $\frac{H_{i}(z^{k})+p_{i}^{k}}{G_{i}(z^{k})+H_{i}(z^{k})+2p_{i}^{k}}\nabla G_{i}(z^{k})+\frac{G_{i}(z^{k})+p_{i}^{k}}{G_{i}(z^{k})+H_{i}(z^{k})+2p_{i}^{k}}\nabla H_{i}(z^{k}).$
As $\epsilon^{k}$ tends to zero, the function $\Phi_{i}^{0}$ is in general not differentiable for $i\in\beta(\bar{z})$. However, if $\Phi_{i}^{0}(z)$ is locally Lipschitz [5, Section 1.2] near $\bar{z}$, the generalized gradient $\partial\Phi_{i}^{0}(\bar{z})$ is generated by a convex hull ([5, Theorem 2.5.1] [6, Eq.(3.1.5)])
$\partial\Phi_{i}^{0}(\bar{z})={\rm conv}\left\{\lim_{s^{K}\to\bar{z}}\nabla\Phi_{i}^{0}(s^{K})\,|\,\nabla\Phi_{i}^{0}(s^{K})\text{ exists}\right\},$
where $\{s^{K}\}$ is any sequence that converges to $\bar{z}$ while avoiding the points where $\Phi_{i}^{0}$ is not differentiable. Noting that $\Phi_{i}^{0}(\bar{z})=\min\{G_{i}(\bar{z}),H_{i}(\bar{z})\}=0$ for $i=1,\dots,m$, we have
$\partial\Phi_{i}^{0}(\bar{z})=\partial\min\{G_{i}(\bar{z}),H_{i}(\bar{z})\}={\rm conv}\{\nabla G_{i}(\bar{z}),\nabla H_{i}(\bar{z})\}.$
For $\delta_{i}\in\partial\Phi_{i}^{0}(\bar{z})$, it follows that ([26, Lemma 1])
$\begin{split}\delta_{i}=\theta_{i}\nabla G_{i}(\bar{z})+(1-\theta_{i})\nabla H_{i}(\bar{z}),&\quad\theta_{i}\in[0,1],\\ \theta_{i}G_{i}(\bar{z})=0,\\ (1-\theta_{i})H_{i}(\bar{z})=0.\end{split}$
Therefore, as $\epsilon^{k}\to 0$, the gradient of $\Phi_{i}^{\epsilon}$ tends to
$\delta_{i}=\left\{\begin{aligned} &\nabla G_{i}(\bar{z}),&&i\in\alpha(\bar{z}),\\ &\nabla H_{i}(\bar{z}),&&i\in\gamma(\bar{z}),\\ &\theta_{i}\nabla G_{i}(\bar{z})+(1-\theta_{i})\nabla H_{i}(\bar{z}),&&i\in\beta(\bar{z}),\end{aligned}\right$. (26)
where $\theta_{i}\in[0,1]$.
Boundedness of multipliers. Without loss of generality, we have the vector of the multipliers $u^{k}\neq 0$ (otherwise $z^{k}$ is an unconstrained local minimum). Let
$\begin{gathered}\Delta^{k}=\sqrt{1+\sum_{i\in I_{g}(z^{k})}(u_{i}^{g,k})^{2}+\sum_{i\in I_{h}(z^{k})}(u_{i}^{h,k})^{2}+\sum_{i=1}^{m}(u^{\Phi,k}_{i})^{2}},\\ \mu^{k}=\frac{1}{\Delta^{k}},\quad\nu_{i}^{g,k}=\frac{u_{i}^{g,k}}{\Delta^{k}},\quad\nu_{i}^{h,k}=\frac{u_{i}^{h,k}}{\Delta^{k}},\quad\nu_{i}^{\Phi,k}=\frac{u^{\Phi,k}_{i}}{\Delta^{k}}.\end{gathered}$ (27)
Dividing (24) by $\Delta^{k}$, we obtain
$0=\mu^{k}\nabla f(z^{k})+\sum_{i\in I_{g}(z^{k})}\nu_{i}^{g,k}\nabla g_{i}(z^{k})+\sum_{i\in I_{h}(z^{k})}\nu_{i}^{h,k}\nabla h_{i}(z^{k})\\ -\sum_{i\in\alpha(\bar{z})}\nu_{i}^{\Phi,k}\nabla\Phi_{i}^{\epsilon}(z^{k})-\sum_{i\in\gamma(\bar{z})}\nu_{i}^{\Phi,k}\nabla\Phi_{i}^{\epsilon}(z^{k})-\sum_{i\in\beta(\bar{z})}\nu_{i}^{\Phi,k}\nabla\Phi_{i}^{\epsilon}(z^{k}).$ (28)
Since we have
$(\mu^{k})^{2}+\sum_{i\in I_{g}(z^{k})}(\nu_{i}^{g,k})^{2}+\sum_{i\in I_{h}(z^{k})}(\nu_{i}^{h,k})^{2}+\sum_{i=1}^{m}(\nu^{\Phi,k}_{i})^{2}=1,$
the sequence $\{(\mu^{k},\nu^{g,k},\nu^{h,k},\nu^{\Phi,k})\}$ is bounded and must converge to some limit $(\bar{\mu},\bar{\nu}^{g},\bar{\nu}^{h},\bar{\nu}^{\Phi})$. It follows from (28) that this limit must satisfy
$0=\bar{\mu}\nabla f(\bar{z})+\sum_{i\in I_{g}(\bar{z})}\bar{\nu}_{i}^{g}\nabla g_{i}(\bar{z})+\sum_{i\in I_{h}(\bar{z})}\bar{\nu}_{i}^{h}\nabla h_{i}(\bar{z})$
$-\sum_{i\in\alpha(\bar{z})}\bar{\nu}^{\Phi}_{i}\nabla G_{i}(\bar{z})-\sum_{i\in\gamma(\bar{z})}\bar{\nu}^{\Phi}_{i}\nabla H_{i}(\bar{z})-\sum_{i\in\beta(\bar{z})}\bar{\nu}^{\Phi}_{i}\left[\theta_{i}\nabla G_{i}(\bar{z})+(1-\theta_{i})\nabla H_{i}(\bar{z})\right],$
where (26) has been used to characterize the derivatives at $\bar{z}$, and $\bar{\mu},\bar{\nu}^{g}\geq 0$ because of (27). Now suppose that $\mu^{k}$ vanishes in the limit, namely, $\bar{\mu}=0$. Then the above equality contradicts the MPCC-MFCQ assumption at $\bar{z}$. Therefore, $\bar{\mu}>0$, and this also implies that the sequence $\{\Delta^{k}\}$ and therefore $\{(u^{g,k},u^{h,k},u^{\Phi,k})\}$ are bounded. This proves the first claim.
Weak and C- stationarity. Without loss of generality, letting $\bar{\mu}=1$ and $\bar{u}=(\bar{u}^{g},\bar{u}^{h},\bar{u}^{\Phi})$ with $\bar{u}^{g}\geq 0$ be the multipliers associated with $\bar{z}$, we obtain
$0=\nabla f(\bar{z})+\sum_{i\in I_{g}(\bar{z})}\bar{u}_{i}^{g}\nabla g_{i}(\bar{z})+\sum_{i\in I_{h}(\bar{z})}\bar{u}_{i}^{h}\nabla h_{i}(\bar{z})$
$-\sum_{i\in\alpha(\bar{z})}\bar{u}^{\Phi}_{i}\nabla G_{i}(\bar{z})-\sum_{i\in\gamma(\bar{z})}\bar{u}^{\Phi}_{i}\nabla H_{i}(\bar{z})-\sum_{i\in\beta(\bar{z})}\bar{u}^{\Phi}_{i}\left[\theta_{i}\nabla G_{i}(\bar{z})+(1-\theta_{i})\nabla H_{i}(\bar{z})\right],$
for some $\theta_{i}\in[0,1]$. Thus $\bar{z}$ satisfies the weak stationarity condition (3), with the MPCC multipliers given by (25). Moreover, $\bar{z}$ is C-stationary because
$\bar{\lambda}^{G}_{i}\cdot\bar{\lambda}^{H}_{i}=(\bar{u}^{\Phi}_{i})^{2}\theta_{i}(1-\theta_{i})\geq 0,\quad\forall i\in\beta(\bar{z}).$ (29)
This proves the second claim. The third claim follows directly from (25). This completes the proof. ∎
3.5 M-stationarity Convergence
The following discussion is inspired by the concept of optimal active set in nonlinear programming and its application to MPCCs in [12]. We introduce two assumptions below.
  1. (A1)
    There exists an optimal active set $\mathcal{A}^{*}$ at $\bar{z}$ that is given by
    $\mathcal{A}^{*}=\{J_{g},I_{h}(\bar{z}),J_{G},J_{H}\}\subseteq\{I_{g}(\bar{z}),I_{h}(\bar{z}),\alpha(\bar{z})\cup\beta(\bar{z}),\gamma(\bar{z})\cup\beta(\bar{z})\},$
    such that
    1. (i)
      $J_{G}\cup J_{H}=\{1,\dots,m\}$;
    2. (ii)
      the gradients $\nabla g_{i}(\bar{z}),\nabla h_{i}(\bar{z}),\nabla G_{i}(\bar{z}),\nabla H_{i}(\bar{z})$ in $\mathcal{A}^{*}$ are linearly independent;
    3. (iii)
      weak stationarity at $\bar{z}$ holds for the set $\mathcal{A}^{*}$, namely, there exist Lagrange multipliers satisfying the weak stationarity condition with $\bar{\lambda}_{i}^{g}\geq 0$ for $i\in J_{g}$, $\bar{\lambda}_{i}^{g}=0$ for $i\notin J_{g}$, $\bar{\lambda}_{i}^{G}=0$ for $i\notin J_{G}$, and $\bar{\lambda}_{i}^{H}=0$ for $i\notin J_{H}$.
  2. (A2)
    At a KKT point $z^{k}$ of BA($\epsilon^{k}$), the reduced Hessian of the Lagrangian (21) is bounded below in the sense that
    $d^{T}\nabla_{zz}\mathcal{L}(z^{k},u^{k})d>-\infty,\quad\forall d\in\mathcal{T}_{\rm BA}^{\rm lin}(z^{k}),$
    where
    $\mathcal{T}_{\rm BA}^{\rm lin}(z^{k})=\{d\,|$ $\nabla g_{i}(z^{k})^{T}d\leq 0,$ $\forall i\in J_{g},$
    $\nabla h_{i}(z^{k})^{T}d=0,$ $\forall i\in I_{h}(z^{k}),$
    $\nabla\Phi_{i}^{\epsilon}(z^{k})^{T}d=0,$ $i=1,\dots,m\}.$
Theorem 3.2.
Suppose that $\bar{z}$ is generated from the sequence described in Theorem 3.1. In addition to the assumptions of Theorem 3.1, suppose that the condition (A1) and MPCC-CRCQ hold at $\bar{z}$, and the condition (A2) holds at every $z^{k}$ when $\epsilon^{k}>0$ sufficiently small. Then $\bar{z}$ is an M-stationary point of MPCC (1).
Proof.
For the purpose of deriving a contradiction, assume that $\bar{z}$ is not M-stationary. Then, by Theorem 3.1, $\bar{z}$ is C-stationary and there exists an index $i_{0}\in\beta(\bar{z})\cap\mathcal{A}^{*}$ such that
$\bar{\lambda}_{i_{0}}^{G}$ $=\bar{u}_{i_{0}}^{\Phi}\theta_{i_{0}}<0,$ (30)
$\bar{\lambda}_{i_{0}}^{H}$ $=\bar{u}_{i_{0}}^{\Phi}(1-\theta_{i_{0}})<0.$
This implies that $\bar{u}_{i_{0}}^{\Phi}<0$ and $0<\theta_{i_{0}}<1$.
Because of the assumption (A1) and MPCC-CRCQ, the gradients in $\mathcal{A}^{*}$ are also linearly independent at $z^{k}$ for $\epsilon^{k}$ sufficiently small. Thus, we can choose a sequence of directions $\{d^{k}\}$ with
$\nabla g_{i}(z^{k})^{T}d^{k}$ $=0,\quad i\in J_{g},$ (31)
$\nabla h_{i}(z^{k})^{T}d^{k}$ $=0,\quad i\in I_{h}(z^{k}),$
$\nabla G_{i}(z^{k})^{T}d^{k}$ $=0,\quad i\in J_{G}\setminus\{i_{0}\},$
$\nabla H_{i}(z^{k})^{T}d^{k}$ $=0,\quad i\in J_{H}\setminus\{i_{0}\},$
$\nabla G_{i_{0}}(z^{k})^{T}d^{k}$ $=\kappa_{G}=\nabla_{H}\Phi^{\epsilon}_{i_{0}}(z^{k}),$
$\nabla H_{i_{0}}(z^{k})^{T}d^{k}$ $=\kappa_{H}=-\nabla_{G}\Phi^{\epsilon}_{i_{0}}(z^{k}).$
The sequence $\{d^{k}\}$ is well-defined, because the coefficient matrix is of full rank and $\kappa_{G},\kappa_{H}$ are bounded (since $\nabla_{G}\Phi_{i_{0}}^{\epsilon}(z^{k})+\nabla_{H}\Phi_{i_{0}}^{\epsilon}(z^{k})=1$ from (23)). Note that $d^{k}\in\mathcal{T}_{\rm BA}^{\rm lin}(z^{k})$. Combining the definition of $d^{k}$ in (31) and the derivatives of $\Phi_{i_{0}}^{\epsilon}$ in (23), we obtain the contribution of the constraint $\Phi_{i_{0}}^{\epsilon}(z^{k})+p_{i_{0}}^{k}=0$ to the reduced Hessian $(d^{k})^{T}\nabla_{zz}\mathcal{L}(z^{k},u^{k})d^{k}$, that is,
- $u_{i_{0}}^{\Phi,k}(d^{k})^{T}\nabla_{zz}\Phi^{\epsilon}_{i_{0}}(z^{k})d^{k}$
$=$ $-u_{i_{0}}^{\Phi,k}(d^{k})^{T}\bigl[\nabla_{G}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla_{zz}G_{i_{0}}(z^{k})+\nabla_{H}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla_{zz}H_{i_{0}}(z^{k})$
$+\nabla_{GG}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla G_{i_{0}}(z^{k})\nabla G_{i_{0}}(z^{k})^{T}+\nabla_{GH}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla G_{i_{0}}(z^{k})\nabla H_{i_{0}}(z^{k})^{T}$
$+\nabla_{HG}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla H_{i_{0}}(z^{k})\nabla G_{i_{0}}(z^{k})^{T}+\nabla_{HH}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla H_{i_{0}}(z^{k})\nabla H_{i_{0}}(z^{k})^{T}\bigr]d^{k}$
$=$ $-u_{i_{0}}^{\Phi,k}(d^{k})^{T}\nabla_{G}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla_{zz}G_{i_{0}}(z^{k})d^{k}-u_{i_{0}}^{\Phi,k}(d^{k})^{T}\nabla_{H}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla_{zz}H_{i_{0}}(z^{k})d^{k}$
$+\frac{2}{G_{i_{0}}(z^{k})+H_{i_{0}}(z^{k})+2p_{i_{0}}^{k}}\nabla_{G}\Phi^{\epsilon}_{i_{0}}(z^{k})\nabla_{H}\Phi^{\epsilon}_{i_{0}}(z^{k})u_{i_{0}}^{\Phi,k}.$
In the last equality, the first two terms are bounded; in the third term,
$\nabla_{G}\Phi_{i_{0}}^{\epsilon}(z^{k}),\nabla_{H}\Phi_{i_{0}}^{\epsilon}(z^{k})>0$
are bounded, $u_{i_{0}}^{\Phi,k}$ tends to $\bar{u}_{i_{0}}^{\Phi}<0$, while $G_{i_{0}}(z^{k})$, $H_{i_{0}}(z^{k})$, and $p_{i_{0}}^{k}$ tend to zero. As a result, we have
$u_{i_{0}}^{\Phi,k}(d^{k})^{T}\nabla_{zz}\Phi^{\epsilon}_{i_{0}}(z^{k})d^{k}\to-\infty.$ (32)
Since all other terms in the reduced Hessian $(d^{k})^{T}\nabla_{zz}\mathcal{L}(z^{k},u^{k})d^{k}$ are bounded, we obtain that (32) contradicts (A2). Hence, our initial hypothesis (30) must be false and $\bar{z}$ is M-stationary. ∎
3.6 Inequality Variant of BA($\epsilon$)
To further explore convergence properties of the Bounding Algorithm, it is beneficial to take advantage of an inequality variant of the problem BA($\epsilon$). We note that this variant is a modification of the Lin-Fukushima algorithm [24], which we call MLF. In this section, a better understanding of the behavior of MLF and BA leads to a simplified LPCC in B-stationarity verification.
Consider an inequality variant of the problem BA($\epsilon$), which is given by
[…]
${\rm s.t.}$ $g(z)\leq 0,$ $u^{g}$
$h(z)=0,$ $u^{h}$
$-\epsilon/2\leq\Phi^{\epsilon}_{i}(z)\leq 0,\;i=1,\dots,m.$ $u^{\Phi}_{L,i},u^{\Phi}_{U,i}$
For a sequence of positive scalars $\epsilon^{k}\to 0$, solving problems MLF($\epsilon^{k}$) generates a sequence $\{z^{k}\}\to\bar{z}$, where every $z^{k}$ is a KKT point of MLF($\epsilon^{k}$). At every point $z^{k}$ we have multipliers $u^{k}=(u^{g,k},u^{h,k},u^{\Phi,k}_{L},u^{\Phi,k}_{U})$ with $u^{g,k}\geq 0$ and $0\leq u_{L,i}^{\Phi,k}\perp u_{U,i}^{\Phi,k}\geq 0$ for $i=1,\dots,m$, such that
$0=\nabla f(z^{k})+\sum_{i\in I_{g}(z^{k})}u_{i}^{g,k}\nabla g_{i}(z^{k})+\sum_{i\in I_{h}(z^{k})}u_{i}^{h,k}\nabla h_{i}(z^{k})-\sum_{i=1}^{m}(u^{\Phi,k}_{L,i}-u^{\Phi,k}_{U,i})\nabla\Phi_{i}^{\epsilon}(z^{k}).$ (34)
Comparing the problem formulations (19) and (33), and the KKT conditions (24) and (34), gives the relations between BA($\epsilon^{k}$) and MLF($\epsilon^{k}$):
$p_{i}^{k}=\epsilon^{k}/2$ $\Leftrightarrow\text{lower bound of }\Phi_{i}^{\epsilon}(z^{k})\text{ is active, and }u_{L,i}^{\Phi,k}\geq 0,$ (35)
$p_{i}^{k}=0$ $\Leftrightarrow\text{upper bound of }\Phi_{i}^{\epsilon}(z^{k})\text{ is active, and }u_{U,i}^{\Phi,k}\geq 0,$
$u^{\Phi,k}$ $=u_{L}^{\Phi,k}-u_{U}^{\Phi,k}.$
In view of the relation between their multipliers, the convergence results established for BA in Theorems 3.1 and 3.2 also hold for MLF, by replacing $\bar{u}^{\Phi}$ with $\bar{u}_{L}^{\Phi}-\bar{u}_{U}^{\Phi}$. In particular, making this substitution in (25) gives the MPCC multipliers at a limit point $\bar{z}$ of MLF:
$\bar{\lambda}^{g}$ $=\bar{u}^{g}=\lim_{k\to\infty}u^{g,k},$ (36)
$\bar{\lambda}^{h}$ $=\bar{u}^{h}=\lim_{k\to\infty}u^{h,k},$
$\bar{\lambda}_{i}^{G}$ $=\left\{\begin{aligned} &\bar{u}^{\Phi}_{L,i}-\bar{u}_{U,i}^{\Phi}=\lim_{k\to\infty}(u_{L,i}^{\Phi,k}-u_{U,i}^{\Phi,k}),&&i\in\alpha(\bar{z})\\ &(\bar{u}^{\Phi}_{L,i}-\bar{u}_{U,i}^{\Phi})\theta_{i},&&i\in\beta(\bar{z}),\end{aligned}\right$.
$\bar{\lambda}_{i}^{H}$ $=\left\{\begin{aligned} &\bar{u}^{\Phi}_{L,i}-\bar{u}_{U,i}^{\Phi}=\lim_{k\to\infty}(u_{L,i}^{\Phi,k}-u_{U,i}^{\Phi,k}),&&i\in\gamma(\bar{z})\\ &(\bar{u}^{\Phi}_{L,i}-\bar{u}_{U,i}^{\Phi})(1-\theta_{i}),&&i\in\beta(\bar{z}).\end{aligned}\right$.
Numerical experience demonstrates that when $\bar{z}$ is not S-stationary, namely, there exists a subset
$\Omega\subseteq\beta(\bar{z}),\text{ such that }\bar{\lambda}_{\Omega}^{G},\bar{\lambda}_{\Omega}^{H}\leq 0,$ (37)
a sequence $\{z^{k}\}$ generated by MLF converges to $\bar{z}$ from the upper bounds of the constraints $\epsilon^{k}/2\leq\Phi_{\Omega}^{\epsilon}(z)\leq 0$, thus $u_{L,\Omega}^{\Phi,k}-u_{U,\Omega}^{\Phi,k}<0$ for all $k$ sufficiently large, and $(\bar{u}_{L,\Omega}^{\Phi}-\bar{u}_{U,\Omega}^{\Phi})<0$ in the limit. In parallel with this observation, a sequence $\{z^{k}\}$ generated by BA converges to $\bar{z}$ with the parameters $p_{\Omega}^{k}=0$ for the constraints $\Phi_{\Omega}^{\epsilon}(z)+p_{\Omega}^{k}=0$, thus the corresponding multipliers $u_{\Omega}^{\Phi,k}<0$ (as implied by (35)), and $\bar{u}_{\Omega}^{\Phi}<0$ in the limit.
Behind these observations lie fundamental reasons that explain the behavior of MLF and BA when approaching a non-strongly stationary point. The following explanation takes advantage of the MLF formulation; but it also addresses the interpretation of BA because of the relation (35) between these two methods.
At a feasible point $z$ of MLF($\epsilon^{k}$), define the index sets
$I^{\Phi}_{L}(z)=\{i\,|\,\Phi_{i}^{\epsilon}(z)$ $=-\epsilon^{k}/2\},$
$I^{\Phi}_{U}(z)=\{i\,|\,\Phi_{i}^{\epsilon}(z)$ $=0\}.$
The constraint $\epsilon^{k}/2\leq\Phi_{i}^{\epsilon}(z)\leq 0$ requires that
$(G_{i}(z)+\tfrac{\epsilon^{k}}{2})(H_{i}(z)+\tfrac{\epsilon^{k}}{2})$ $\geq(\tfrac{\epsilon^{k}}{2})^{2},$
$G_{i}(z)H_{i}(z)$ $\leq(\tfrac{\epsilon^{k}}{2})^{2},$
and at the lower and upper bounds we have
$G_{i}(z)+\tfrac{\epsilon^{k}}{2}>0,\,H_{i}(z)+\tfrac{\epsilon^{k}}{2}>0,\,(G_{i}(z)+\tfrac{\epsilon^{k}}{2})(H_{i}(z)+\tfrac{\epsilon^{k}}{2})=(\tfrac{\epsilon^{k}}{2})^{2},$ $\forall i\in I^{\Phi}_{L}(z),$
$G_{i}(z)>0,\,H_{i}(z)>0,\,G_{i}(z)H_{i}(z)=(\tfrac{\epsilon^{k}}{2})^{2},$ $\forall i\in I^{\Phi}_{U}(z).$
Therefore, the feasible region of MLF($\epsilon^{k}$) includes the feasible region of MPCC (1); also, it restricts the feasible region of RNLP (6) from above by enforcing $\Phi_{i}^{\epsilon}(z)\leq 0$, and extends the region a little below by using the relaxed lower bounds $\Phi_{i}^{\epsilon}(z)\geq-\epsilon^{k}/2$ to allow for small perturbations $G_{i}(z)<0$ or $H_{i}(z)<0$. Suppose that there exists a subset $\Omega\subseteq\{1,\dots,m\}$, such that RNLP (6) is minimized at $G_{\Omega}(z)>0$ and $H_{\Omega}(z)>0$. As the solutions of the RNLP locate outside of the feasible region of the MPCC, no local minimizer of the MPCC can be S-stationary. In such circumstance, the RNLP constrained additionally by $\Phi_{\Omega}^{\epsilon}(z)\leq 0$ achieves the minimal cost on the boundaries of $\Phi_{\Omega}^{\epsilon}(z)\leq 0$ for every $\epsilon^{k}>0$. For MLF($\epsilon^{k}$), it may have the same minimizer as the additionally constrained RNLP, or have a better solution on the lower bound of $\Phi_{i}^{\epsilon}(z)$ for some $i\in\Omega$ and every $\epsilon^{k}>0$ suitably small. In the latter case, for those $i\in\Omega$, we have that in the limit $\bar{u}_{L,i}^{\Phi}-\bar{u}_{U,i}^{\Phi}\geq 0$ and therefore $\bar{\lambda}_{i}^{G},\bar{\lambda}_{i}^{H}\geq 0$ (as indicated by (36)), which contradicts the assumption that RNLP is minimized at $G_{i}(z),H_{i}(z)>0$. Therefore, for every $\epsilon^{k}>0$ suitably small, a local minimizer of MLF($\epsilon^{k}$) is also a local minimizer of the RNLP constrained additionally by $\Phi_{\Omega}^{\epsilon}(z)\leq 0$. This gives rise to the phenomenon that the upper bounds of the constraints $\epsilon^{k}/2\leq\Phi_{\Omega}^{\epsilon}(z)\leq 0$ are active at every $z^{k}$ as $\epsilon^{k}\to 0$. Moreover, we have $\Omega\subseteq\beta(\bar{z})$ because the constantly active upper bounds as $\epsilon^{k}\to 0$ means that $G_{\Omega}(z^{k})>0,H_{\Omega}(z^{k})>0$, and $G_{\Omega}(z^{k})H_{\Omega}(z^{k})=(\epsilon^{k}/2)^{2}$ (componentwise product) for infinitely many $k$.
The above interpretation of the algorithms behavior may help to reduce the cost in verifying whether a limit point is B-stationary. Consider a limit point $\bar{z}$ of BA or MLF, at which there exists a subset
$\Omega\subseteq\beta(\bar{z}),\text{ such that }\bar{u}_{\Omega}^{\Phi}<0\text{ (BA) and }\bar{u}_{L,\Omega}^{\Phi}-\bar{u}_{U,\Omega}^{\Phi}<0\text{ (MLF)}.$
According to (25) and (36), the MPCC multipliers have the property (37). Theorem 2.3 states that B-stationarity can be obtained from piecewise M-stationarity. Since the above discussion has shown that such subset $\Omega$ usually signifies the absence of a S-stationary solution, piecewise M-stationarity can be satisfied by the multipliers (37) only if, for every partition $(\beta_{1},\beta_{2})\in\mathcal{P}(\beta(\bar{z}))$, there are multipliers satisfying
$\bar{\lambda}_{i}^{G}<0,\;\bar{\lambda}_{i}^{H}=0,$ $\forall i\in\beta_{1}\cap\Omega,$ (38)
$\bar{\lambda}_{i}^{G}=0,\;\bar{\lambda}_{i}^{H}<0,$ $\forall i\in\beta_{2}\cap\Omega.$
[…]
$\min$ ${\rm obj}(d)=\nabla f(\bar{z})^{T}d$ (39)
${\rm s.t.}$ $\nabla g_{I}(\bar{z})^{T}d\leq 0,$
$\nabla h(\bar{z})^{T}d=0,$
$\nabla G_{\alpha}(\bar{z})^{T}d=0,$
$\nabla H_{\gamma}(\bar{z})^{T}d=0,$
$\nabla G_{\beta_{1}}(\bar{z})^{T}d=0,$ $\nabla H_{\beta_{1}\setminus\Omega}(\bar{z})^{T}d\geq 0,$
$\nabla G_{\beta_{2}\setminus\Omega}(\bar{z})^{T}d\geq 0,$ $\nabla H_{\beta_{2}}(\bar{z})^{T}d=0.$
In (39), the constraints corresponding to the subset $\Omega$ are excluded from the inequality constraints, because (38) implies that the constraints corresponding to $\bar{\lambda}_{i}^{H}$ for all $i\in\beta_{1}\cap\Omega$, and corresponding to $\bar{\lambda}_{i}^{G}$ for all $i\in\beta_{2}\cap\Omega$, must be locally inactive. As a result, only the constraints indicated in (39) need to be considered to check whether these LPs are solved by $d=0$ or not. This can lead to dealing with fewer inequalities than in (10).
4 Practical Issues
We take a closer look at the behavior of the NCP-based bounding methods (BA and MLF) and the typical regularization scheme proposed in [27], when converging to a limit point $\bar{z}$ that is not S-stationary. This reveals an advantage of the complementarity reformulations based on the smoothed NCP function.
4.1 Unbounded NLP Multipliers and Inaccurate Solution
In the course of seeking a solution of an MPCC, NLP subproblems may encounter unbounded multipliers when approaching a limit point that is not S-stationary. Our numerical experience to date indicates that the NCP-based reformulations BA($\epsilon$) and MLF($\epsilon$) avoid unbounded NLP multipliers (even if MPCC-MFCQ fails at a limit point). The following confirms this observation, by comparing these two methods with the typical regularization scheme proposed in [27]:
[…]
${\rm s.t.}$ $g(z)\leq 0,$ $v^{g}$
$h(z)=0,$ $v^{h}$
$G(z)\geq 0,$ $v^{G}$
$H(z)\geq 0,$ $v^{H}$
$G_{i}(z)H_{i}(z)\leq\epsilon,\;i=1,\dots,m.$ $v^{REG}_{i}$
Solving a sequence of problems REG($\epsilon^{k}$) with the positive scalars $\epsilon^{k}\to 0$, generates a sequence $\{z^{k}\}\to\bar{z}$. Based on stationarity of $z^{k}$ for REG($\epsilon^{k}$), namely,
$0=$ $\nabla f(z^{k})+\sum_{i\in I_{g}(z^{k})}v_{i}^{g,k}\nabla g_{i}(z^{k})+\sum_{i\in I_{h}(z^{k})}v_{i}^{h,k}\nabla h_{i}(z^{k})$
$-\sum_{i=1}^{m}v^{G,k}_{i}\nabla G_{i}(z^{k})-\sum_{i=1}^{m}v^{H,k}_{i}\nabla H_{i}(z^{k})$
$+\sum_{i=1}^{m}v^{REG,k}_{i}\left[H_{i}(z^{k})\nabla G_{i}(z^{k})+G_{i}(z^{k})\nabla H_{i}(z^{k})\right]$
with $v^{g,k},v^{G,k},v^{H,k},v^{REG,k}\geq 0$, the relations between the NLP multipliers $v^{k}=(v^{g,k},v^{h,k},v^{G,k},v^{H,k},v^{REG,k})$ at $z^{k}$ and the MPCC multipliers $\bar{\lambda}=(\bar{\lambda}^{g},\bar{\lambda}^{h},\bar{\lambda}^{G},\bar{\lambda}^{H})$ at $\bar{z}$ can be expressed by (see also [27, Eq.(6) and Theorem 3.1])
$\bar{\lambda}^{g}$ $=\bar{v}^{g}=\lim_{k\to\infty}v^{g,k},$ (40)
$\bar{\lambda}^{h}$ $=\bar{v}^{h}=\lim_{k\to\infty}v^{h,k},$
$\bar{\lambda}_{i}^{G}$ $=\lim_{k\to\infty}\left[v_{i}^{G,k}-v_{i}^{REG,k}H_{i}(z^{k})\right],\;i=1,\dots,m,$
$\bar{\lambda}_{i}^{H}$ $=\lim_{k\to\infty}\left[v_{i}^{H,k}-v_{i}^{REG,k}G_{i}(z^{k})\right],\;i=1,\dots,m.$
It has been proved that $\bar{z}$ is S-stationary for MPCC (1) if and only if it is a stationary point of REG(0) [10, Proposition 4.1], provided that bounded multipliers exist.
Consider the cases where $\bar{z}$ is not S-stationary. In the case where $\bar{z}$ is no better than C-stationary, there exist indices $i\in\beta(\bar{z})$ such that $\bar{\lambda}_{i}^{G}<0,\bar{\lambda}_{i}^{H}<0$. According to (40), the NLP multipliers $v_{i}^{G,k}$ and $v_{i}^{H,k}$ will have a tendency to be less than zero for $k$ sufficiently large, which are not allowed in REG($\epsilon^{k}$). Since
$\lim_{k\to\infty}v_{i}^{G,k}$ $=\bar{\lambda}_{i}^{G}+\lim_{k\to\infty}v_{i}^{REG,k}H_{i}(z^{k}),$ (41)
$\lim_{k\to\infty}v_{i}^{H,k}$ $=\bar{\lambda}_{i}^{H}+\lim_{k\to\infty}v_{i}^{REG,k}G_{i}(z^{k}),$
the multipliers $v_{i}^{REG,k}$ must become very large to ensure $v_{i}^{G,k}$ and $v_{i}^{H,k}$ remain nonnegative. At the same time, $G_{i}(z^{k})$ and $H_{i}(z^{k})$ are prevented from being very close to zero, otherwise $v_{i}^{REG,k}G_{i}(z^{k})$ and $v_{i}^{REG,k}H_{i}(z^{k})$ would be ineffective. As a consequence, it can be observed for $k$ sufficiently large that $v_{i}^{G,k}=0,v_{i}^{H,k}=0,v_{i}^{REG,k}\to\infty$, and $G_{i}(z^{k})$ and $H_{i}(z^{k})$ cannot converge accurately to zero.
In the case that $\bar{z}$ is no better than M-stationary, there exist indices $i\in\beta(\bar{z})$ such that $\bar{\lambda}_{i}^{G}=0,\bar{\lambda}_{i}^{H}<0$ (or the reverse). The relations (40) imply that for $k$ sufficiently large $v_{i}^{H,k}$ has a tendency to be less than zero, which is not a suitable NLP multiplier. We also use (41) to predict the behavior of the REG method. In order to enforce $v_{i}^{H,k}$ nonnegative, the multipliers $v_{i}^{REG,k}$ become very large, and at the same time, $G_{i}(z^{k})$ cannot be very close to zero. The components $H_{i}(z^{k})$ cannot approach zero quickly either, because the constraints $G_{i}(z^{k})H_{i}(z^{k})\leq\epsilon^{k}$ must be kept active for every $\epsilon^{k}>0$. As a result, the observation for $k$ sufficiently large would be the same as the above C-stationary case.
On the other hand, the multipliers for the problems BA($\epsilon^{k}$) and MLF($\epsilon^{k}$) do not have this difficulty. As indicated by the relations (25) and (36), there is no contradiction between the signs of the MPCC multipliers $\bar{\lambda}_{i}^{G},\bar{\lambda}_{i}^{H}$ and of the NLP multipliers $u_{i}^{\Phi,k}$ and $u_{L,i}^{\Phi,k}-u_{U,i}^{\Phi,k}$. In addition, the underlying relation
$\bar{\lambda}_{i}^{G}+\bar{\lambda}_{i}^{H}=\bar{u}_{i}^{\Phi}=\bar{u}_{L,i}^{\Phi}-\bar{u}_{U,i}^{\Phi},\quad\forall i\in\beta(\bar{z})$ (42)
indicates that the NLP multipliers exist whenever the MPCC multipliers do. Therefore, whether $\bar{z}$ is S-stationary or not has little influence on the performance of BA and MLF methods, which is an important difference from the REG method.
4.2 Examples: MPCC Methods Comparison
The following examples illustrate the difference in behavior between the NCP-based bounding methods and the REG regularization method when approaching a non-strongly stationary solution.
As noted earlier, the problem scholtes4 has a B-stationary point that is not strongly stationary. Another example is problem ex9.2.2 from the MacMPEC collection [22], which is given by
[…]
$-x+y\leq 0,$ $\lambda_{1}$
[…]
$-y+s_{2}=0,$ $\lambda_{3}$
[…]
$0\leq s_{i}\perp l_{i}\geq 0,\;i=1,2,3.$ $\sigma^{si},\sigma^{li}$
A local minimizer is $\bar{z}=(\bar{x},\bar{y},\bar{s},\bar{l})$ with
$\bar{x}=10,\;\bar{y}=10,\;\bar{s}=(0,10,10),\;\bar{l}=(0,0,0).$
The weak stationarity conditions at $\bar{z}$ require that
$2\bar{x}-\lambda_{1}+\lambda_{2}+2\lambda_{5}$ $=0,$ (43)
[…]
$\lambda_{2}-\sigma^{s1}$ $=0,$
$\lambda_{3}$ $=0,$
$\lambda_{4}$ $=0,$
$\lambda_{5}-\sigma^{l1}$ $=0,$
$-\lambda_{5}-\sigma^{l2}$ $=0,$
$\lambda_{5}-\sigma^{l3}$ $=0,$
which implies
[…]
$\sigma^{l1}$ $=\lambda_{5},$
and therefore, $\sigma^{s1},\sigma^{l1}$ cannot be both nonnegative. Let $\sigma^{s1}$ or $\sigma^{l1}$ be zero, then we obtain multipliers with $(\sigma^{s1},\sigma^{l1})=(0,-10/3)$ or $(\sigma^{s1},\sigma^{l1})=(-10,0)$, indicating that $\bar{z}$ is piecewise M-stationary.
Numerical results of examples scholtes4 and ex9.2.2 are presented in Tables 1 and 2, respectively. Note that we can produce MPCC-MFCQ failure in these examples simply by introducing an additional constraint, for instance, $z_{1}\geq 0$ or $z_{1}\leq 0$ (to scholtes4), and $s_{1}\geq 0$ or $s_{1}\leq 0$ (to ex9.2.2); for the resulting problems we still get the same results as in these tables. The results indicate that REG method gives rise to large NLP multipliers for the constraints corresponding to the biactive complementary components, and the multipliers get even larger when the regularization parameter $\epsilon$ becomes smaller. At the same time, the convergence is slow and inaccurate, compared to the magnitude of $\epsilon$. In contrast, the multipliers of the NCP-based bounding methods are well behaved. According to (42), their multipliers can be used to derive the MPCC multipliers at a limit point and vice versa. In addition, the accuracy of their solutions (the problem variables and multipliers) is comparable to $\epsilon$.
Table 1: Results of problem scholtes4.
$\epsilon$ scholtes4 BA MLF REG
$p$ $u^{\Phi}$ $u_{L}^{\Phi}$ $u_{U}^{\Phi}$ $v^{z_{1}}$ $v^{z_{2}}$ $v^{REG}$
0 -2 0 2 0 0 1.00E+3
$10^{-6}$ $z_{1}$ 5E-7 5E-7 0.001000
$z_{2}$ 5E-7 5E-7 0.001000
$z_{3}$ 2E-6 2E-6 0.003999
$p$ $u^{\Phi}$ $u_{L}^{\Phi}$ $u_{U}^{\Phi}$ $v^{z_{1}}$ $v^{z_{2}}$ $v^{REG}$
0 -2 0 2 0 0 2.69E+4
$10^{-9}$ $z_{1}$ 5E-10 5E-10 0.000037
$z_{2}$ 5E-10 5E-10 0.000037
$z_{3}$ 2E-9 2E-9 0.000149
$p$ $u^{\Phi}$ $u_{L}^{\Phi}$ $u_{U}^{\Phi}$ $v^{z_{1}}$ $v^{z_{2}}$ $v^{REG}$
0 -2 0 2 0 0 5.02E+4
$10^{-12}$ $z_{1}$ 5E-11 5E-11 0.000020
$z_{2}$ 5E-11 5E-11 0.000020
$z_{3}$ 2E-10 2E-10 0.000080
Table 2: Results of problem ex9.2.2.
$\epsilon$ ex9.2.2 BA MLF REG
$p$ $u^{\Phi}$ $u_{L}^{\Phi}$ $u_{U}^{\Phi}$ $v^{s_{1}}$ $v^{l_{1}}$ $v^{REG}$
0 -5.74 0 5.74 0 0 2.89E+3
$10^{-6}$ $s_{1}$ 3.8E-7 3.8E-7 0.000577
$l_{1}$ 6.5E-7 6.5E-7 0.001732
$p$ $u^{\Phi}$ $u_{L}^{\Phi}$ $u_{U}^{\Phi}$ $v^{s_{1}}$ $v^{l_{1}}$ $v^{REG}$
0 -4.78 0 5.63 0 0 7.85E+4
$10^{-9}$ $s_{1}$ 2.04E-10 3.65E-10 0.000021
$l_{1}$ 1.11E-10 5.96E-10 0.000064
$p$ $u^{\Phi}$ $u_{L}^{\Phi}$ $u_{U}^{\Phi}$ $v^{s_{1}}$ $v^{l_{1}}$ $v^{REG}$
0 -9.94 0 3.34 0 0 1.46E+5
$10^{-12}$ $s_{1}$ 2.94E-11 2.03E-11 0.000011
$l_{1}$ 3.81E-11 1.09E-11 0.000034
More numerical results of the methods BA, MLF, and REG can be found in [31]. In that study, we considered a selection of problems from the MacMPEC collection [22], which have solutions with biactive complementarity constraints, as well as seven MPCC problems drawn from distillation models with up to 1264 variables and 48 complementarity constraints. Based on our numerical experience so far, we summarize features of these MPCC methods as follows. On the other hand, a thorough numerical study is beyond the current scope and a topic for future work.
  • REG: When approaching a S-stationary solution, REG converges quickly and accurately; the accuracy of the solution does not depend on the accuracy of $\epsilon$, in other words, an accurate solution can be obtained even if $\epsilon$ is not so small. On the other hand, when approaching a non-strongly stationary solution, the convergence can be slow and inaccurate, compared to the magnitude of $\epsilon$; unbounded NLP multipliers usually arise.
  • BA/MLF: Whether it is approaching a S-stationary or a non-strongly stationary solution, performance of the method does not vary much. In particular, BA/MLF is not as efficient as REG in the former case; but in the latter case, it is still able to obtain $O(\epsilon)$-accurate solutions and has well-behaved NLP multipliers .
  • To see the influence of singularity and regularization (because of $\Phi^{\epsilon}$) on BA/MLF, we repeat the loop for solving MPCCs to a very small $\epsilon$ and use a solver, IPOPT [30], based on a second-order algorithm. In the case of approaching a S-stationary point, the regularization of the Lagrangian Hessian is mild and does not happen often, which may attribute to the positive parameter $p>0$ (see the Hessian expression (23), and note that in S-stationarity convergence $G_{i}(z^{k})$ and $H_{i}(z^{k})$ tend to zero quickly for every $i\in\beta(\bar{z})$ and the parameter $p_{i}^{k}>0$ as implied by (35)). In the case of approaching a non-strongly stationary point, the Hessian regularization is rarely observed. The reason is that, as we discussed in Section 3.6, the sequence $\{z^{k}\}$ converges to $\bar{z}$ with $G_{i}(z^{k}),H_{i}(z^{k})>0$ for those indices $i$ such that RNLP (6) approaches its minimum with $G_{i}(z^{k})>0$ and $H_{i}(z^{k})>0$, and these positive $G_{i}(z^{k})$ and $H_{i}(z^{k})$ do not cause numerical difficulties.